40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 39 of 337  Not logged in ELOG logo
ID Date Author Type Category Subject
  15051   Wed Nov 27 12:16:52 2019 gautamUpdateLSCITMX and ITMY OSEMs with low and high circulating power

Summary:

The ITMX OSEMs report elevated noise in the 10-100 Hz band when we have high circulating power in the arm cavities, see Attachment #1. Since there is no LSC actuation on the ITMs in this state, this could be a radiation presssure effect, or could be scattered 1064nm light entering the OSEMs. The Oplevs don't report any elevated noise however. ITMY has the OSEM whitening broken for two channels, but the other two channels don't report as significant an increase as ITMX, see Attachment #2. I can't find the status of which OSEMs have the 1064nm blocking filters installed. The local damping loops are rolled off by ~100dB at 30 Hz, so the sensing noise re-injection should be attenuated by this factor, so maybe the OSEM sensor noise isn't the likely culprit. But radiation pressure didn't worsen the length noise in the past, even after our mirror cleaning and the increased PRG.

Quote:

...maybe the opto-mechanical CARM plant is changing as a function of the CARM offset...

Attachment 1: ITMXshadowSensors.pdf
ITMXshadowSensors.pdf
Attachment 2: ITMYshadowSensors.pdf
ITMYshadowSensors.pdf
  15050   Tue Nov 26 18:16:08 2019 ranaUpdateLSCPOX / POY calibration

...maybe the opto-mechanical CARM plant is changing as a function of the CARM offset...

Quote:

Even assuming 50% error in the calibration factors, it's hard to explain the swing of TRX/TRY when the CARM offset is brought to zero.

  • The increase in (admittedly in-loop) CARM noise as the offset is reduced still seems to me to be correlated with the buildup of IR power in the arm cavities.
  15049   Tue Nov 26 17:07:41 2019 gautamUpdateLSCPOX / POY calibration

Summary:

Since we are using the POX and POY photodiodes as out-of-loop sensors for measuring the ALS noise, I decided to double-check their calibrations. I determined the following numbers (for the single arm lock):

POX_I [with 30dB whitening gain]: (8 +/- 1)e-13 m/ct

POY_I [with 18dB whitening gain]: (0.9 +/- 0.1)e-13 m/ct

With this calibration, I measured the in-loop spectra of the XARM and YARM error-points when they are locked - they line up well, see Attachment #1. Note that these numbers are close to what we determined some time ago using the same method (I drove the ITMs then, but yesterday I drove the ETMs, so maybe the more accurate measure of uncertainty is the difference between the two measurements).

Attachment #2 shows the out-of-loop spectra sensed by these photodiodes with this calibration applied, when the arms are under control using ALS beat frequencies as the error signals, and controlled in the CARM/DARM basis. Need to think about why there is such a difference between the two signals.

Methodology:

The procedure used was the same as that outlined here.

  • I started by calibrating the AS55_Q output with the free-swinging Michelson.
  • Next, I lock the Michelson and calibrate the BS and ITM actuators using the newly calibrated AS55_Q.
  • Next, I calibrate the ETM actuator gains by measuring the ratio of response in POX/POY of driving the (unknown) ETMs and the (known) ITMs.
  • Finally, I calibrate the POX/POY photodiodes by driving the ETMs by a known amount of meters (at ~310 Hz where the loop gain is negligible because of the sensing matrix measurement notches).

Summary of DC actuator gains:

Optic Series resistance [ohms] x3 Analog gain? x3 Digital gain? DC gain [nm/ct]
BS 100 No Yes 9.48 +/- 0.01
ITMX 400 No Yes 2.42 +/- 0.01
ITMY 400 No Yes 2.41 +/- 0.01
ETMX 2.2k Yes No 1.23 +/- 0.02
ETMY 400 Yes No 6.62 +/- 0.12

The quoted values of the DC gain are for counts seen at the output of the LSC filter bank. I've attempted to show that once we account for the different series resistance and some extra gains between the output of the LSC filter bank and the actual coil, things are fairly consistent.

Some remarks:

  • I do not understand why we need an extra 12dB of whitening gain on the POX channel to get similar PDH fringe height as the POY channel. The light level on these photodiodes is the same, and the RF transimpedances at 11 MHz are also close according to the wiki (3kohm for POX, 2kohm for POY).
  • At night-time, the ALS noise did indeed get reduced compared to what I measured earlier in the evening.
  • Even assuming 50% error in the calibration factors, it's hard to explain the swing of TRX/TRY when the CARM offset is brought to zero.
  • The increase in (admittedly in-loop) CARM noise as the offset is reduced still seems to me to be correlated with the buildup of IR power in the arm cavities.
Attachment 1: POX_POY_sensorNoise.pdf
POX_POY_sensorNoise.pdf
Attachment 2: ALSnoise_20191125.pdf
ALSnoise_20191125.pdf
  15048   Tue Nov 26 13:33:33 2019 YehonathanUpdateCamerasMC2 Camera rotated by 90 degrees

MC2 analog camera was rotated by 90 degrees. Orientation correctness was verified by exciting the MC2 Yaw degree of freedom.

Attached before and after photos of the camera setup.

Attachment 1: MC2AnalogCameraAfter.jpg
MC2AnalogCameraAfter.jpg
Attachment 2: MC2AnalogCameraBefore.jpg
MC2AnalogCameraBefore.jpg
  15047   Mon Nov 25 22:10:26 2019 shrutiUpdateNoiseBudgetDiagnostics

This is to help troubleshoot the excess noise measured earlier.

The following channels were measured at GPS times 1258586880 s and 1258597457 s, corresponding to low and high Power Recycling Gain (PRG) respectively.

Excess noise was seen between 25-110 Hz in the high PRG case when compared to the low PRG case in the following channels:

 

C1:LSC-CARM-IN1_DQ (shown in Attachment 1 where the reference is low PRG)

C1:ALS-Y_ERR_MON_OUT_DQ

C1:ALS-BEAT{X,Y}_FINE_PHASE_OUT_DQ

C1:SUS-ETM{X,Y} _SENSOR_{LL,LR,UL,UR}

C1:ALS-TRX_OUT_DQ

 

Surprisingly, it was also seen to a smaller extent in (refer Attachment 3)

C1:SUS-ITMX_SENSOR_{LL,LR,UL,UR} 

 

A different type of noise spectrum, attributed to known electronic effects, was observed for

C1:SUS-ITMY_SENSOR_{LL,UL}    (refer Attachment 2)

 

These did not show any significant change in the noise spectrum:

C1:LSC-DARM-IN1_DQ (shown in Attachment 1 where the reference is low PRG)

C1:ALS-X_ERR_MON_OUT_DQ

C1:ALS-TRY_OUT_DQ

C1:SUS-ITMY_SENSOR_{LL,LR,UL,UR} 

C1:SUS-ITMY_SENSOR_{LR,UR}  (refer Attachment 2)

 

Broadband noise in:

C1:LSC-PO{X,Y}11_I_ERR_DQ

 

Attachment 1: LSC.pdf
LSC.pdf
Attachment 2: ITMY.pdf
ITMY.pdf
Attachment 3: ITMX_L.pdf
ITMX_L.pdf
  15046   Mon Nov 25 19:11:22 2019 gautamUpdateLSCALS noise re-look

I re-checked the ALS noise in the following configurations:

  • PRM is misaligned.
  • Michelson is not locked.
  • TRX/TRY is maintained at ~1.
  1. Arm lengths are controlled using POX/POY as a sensor, and the ETMs as actuators [orange traces in Attachment #1].
    • EX laser frequency is locked to the arm cavity length using the end PDH servo.
    • ALS beat note frequency fluctuations are read out using the calibrated DFD channels.
    • In this config, the DFD outputs are the out-of-loop sensor.
  2. Arm lengths are controlled using the ALS beat frequencies as a sensors [blue traces in Attachment #1]
    • The control is no longer in the XARM/YARM basis, but in the CARM/DARM basis.
    • The CARM actuator is MC2, the DARM actuator is an admixture of the ETMs (equal magnitude of output matrix element, opposite sign).
    • The calibrated POX/POY photodiodes are used as the out-of-loop sensor in this config.

The RMS noise sensed by POX/POY is ~20pm, which is somewhat higher than the best I've seen (maybe the arms are moving more at the time of measurement or the AUX PDH loops need a bit of touching up). But the orange traces in the top row of Attachment #1 are already ~x2 better than the equivalent traces from when we were using the green beams to make the beats. So it's hard to explain the 0-300 fluctuations in the arm powers when the CARM offset is reduced to 0 - i.e. the ALS noise is becoming worse as I reduce the CARM offset (= have more circulating power compared to the conditions of this test). I assume the transmission is Lorentzian, in which case even if we have 5x the CARM linewidth worth of ALS noise, we should see the arm power fluctuate between ~10 and 300. 

* I notice that a big jump in the RMS sensed by POX/POY comes from the 24 Hz peak, which is presumably the Roll mode coupling to length - maybe a ResG can make the situation better. The high frequency noise can also be probably rolled off better.

Attachment 1: ALSnoise.pdf
ALSnoise.pdf
  15045   Fri Nov 22 00:54:14 2019 gautamUpdateLSClocking notes

[KA, GV]

There was no shaking (that disturbed the locking) tonight!

  1. REFL165 Demod phase was adjusted from -111deg to -125deg. To minimize coherence b/w MICH and PRCL.
  2. MICH 3f loop gain changed to 0.3.
  3. If the POP mode shape looks weird, it probably means that the PRM is sligntly misaligned. Tweaking the alignment improves PRMI stability and also makes the arm buildup higher.
  4. Ditto for MICH - slightly touching up the BS alignment can lower ASDC.
  5. Main finding tonight was that the ALS noise seems to get degraded as a function of the CARM offset! As a result of this, CARM goes through several linewidths, and the arm transmission fluctuates wildly.
    • We suspect some scattered light shenanigans. It is not clear to me why this is happening. Possibilities:
    • Scattered ETM transmission somehow makes it into the fiber coupler and degrades the ALS noise.
    • Sacttered ETM transmission makes it onto the Green PDH photodiode and degrades the ALS noise.
    • Backscatter into the PSL degrades the ALS noise.
    • Shadow sensors of either the ITMs, ETMs, BS, or PRM don't have 1064nm filters and get scatterd light, making the cavity length noise worse.
    • Other possibilities?

The problem is hard to debug because we are feeding back on the ETMs, BS and PRM, and at the low CARM offset (= high PRG), all the DoFs are cross coupled strongly so just by looking at error/control signals, I can't directly determine where the noise is originating. The fact that the ALS CARM spectrum shows a noise increase suggests that the problem has to do with the test masses, PSL, IMC, or end green PDH setups.

My plan is to do a systematic campaign and eliminate some of these possibilities - e.g. install some baffling around the fiber coupler and the end green PDH photodiodes and see if there is any improvement in the situation.

* In attachment #1, the "Ref" traces are when the CARM offset is 0, and the arms are buzzing in and out of resonance. The non-reference traces are for when the CARM offset is ~28kHz (so several linewidths away from resonance).

Attachment 1: ALSnoiseIncrease.pdf
ALSnoiseIncrease.pdf
  15044   Thu Nov 21 19:08:58 2019 gautamUpdateLSCHigh BW lock of Y arm length to PSL frequency

Summary:

The Y arm cavity length was locked to the PSL frequency with ~26kHz UGF, and 25 degrees phase margin. Slow actuation was done on ETMY using CM_Slow as an error signal, while fast actuation was done on the IMC error point via the IN2 input of the IMC servo board. Attachment #1 shows the comparison of the in-loop error signal spectra with only slow actuation and with the full CM loop engaged.

Details:

  1. LSC enable OFF.
  2. Configure the CM board for locking:
    • CM board IN1 gain = 25dB.
    • CM_Slow whitening gain = +18dB, make sure the offsets are correctly set. CM_Slow filter bank = -0.015.
    • CM_Slow-->YARM matrix element in LSC input matrix is -2.5.
    • YARM-->ETMY matrix element in LSC Output matrix is 1.
    • AO gain set to +5dB. IMC Servo board IN2 gain starts at -32dB, the path is disabled. The polarity is Plus.
    • Usual YARM FM triggers are set (FM1, FM2, FM3, FM6, FM8), usual YARM servo gain is used (0.01), usual triggering conditions (ON @ TRY>0.3, OFF @ TRY < 0.1), usual power normalization by TRY.
  3. Enable LSC mode, wait for the arm to acquire lock.
  4. Once the digital boosts are engaged, enable the IMC IN2 path, ramp up the gain to -2 dB. Note that this IN2 path is AC coupled, according to this elog. The corner frequency is 1/2/pi/2e3/6.8uF ~11 Hz. This was confirmed by measurement, see Attachment #3. I couldn't find a 2-pin LEMO-->BNC adaptor so I measured at the BNC connector for the IN2 input, which according to the schematic is shorted to the LEMO (which is what we use for the AO path).
  5. Enable the CM board boost.
  6. Ramp up the CM board IN1 gain to +31dB. In this config, the CM_Slow signal is ~18,000 cts pk (with the +18dB whitening gain), so not saturating the ADC.
  7. Ramp up the IMC IN2 gain to 3dB, engage 2 Super Boosts (can't turn on the third). Limiter is always ON.
  8. Use the CM board error point offset adjust to zero the POY11_I error signal average value - there seems to be some offsets when engaging the boosts. The value I used was 0.9 V (this is internally divided by 40 on the CM board).
  9. Whiten the CM_Slow signal - this doesn't seem to have any impact on the noise anywhere.

I hypothesize that the high-frequency noise (>100 Hz) is higher for POY than POX in Attachment #1 because I am using the "MON" port of the demod board - this has a gain of 2, and there could also be some flaky components in this path, hence the high frequency noise is a factor of a few greater in the POY spectrum relative to the POX spectrum (which is using the main demodulated output). For REFL11, we have a low noise preamp generating the input signal so I don't think we need to worry about this too much.

The PC Drive RMS didn't look any stranger than it usually does for the duration of the lock.

Attachment #2 shows the OLTF of the locking servo with the final gains / settings, which are in bold. The loop is maybe a bit marginal, could possibly benefit from backing off one of the super boosts. But the arm has stayed locked for >1 hour.

The purpose of this test was to verify the functionality of the CM board and also the IN2 of the IMC servo board in a low-pressure environment. Once I confirm that the modelled OLTF lines up with the measured, I will call this test a success, and move on to looking at REFL11 in the arms on ALS, PRMI on 3f config. I am returning the REFL11 signal to the input of the CM board, but the SR785 remains hooked up.

Unrelated to this work - PMC alignment was tweaked to improve input power to IMC by ~5%.

Attachment 1: highBW_POY.pdf
highBW_POY.pdf
Attachment 2: CM_UGF.pdf
CM_UGF.pdf
Attachment 3: IN2_ACcoupling.pdf
IN2_ACcoupling.pdf
  15043   Thu Nov 21 13:14:33 2019 KojiUpdateLSCCM board study

One of the differences between the direct POY and the CM_SLOW POY is the presence of the CM Servo gain stages. So this might mean that you need to move some of the whitening gain to the CM IN1 gain.

  15042   Thu Nov 21 12:46:22 2019 gautamUpdateLSCCM board study

In preparation for trying out some high-bandwidth Y arm cavity locking using the CM board, I hooked up the POY11_Q_Mon channel of the POY11 demod board to the IN1 of the CM board (and disconnected the usual REFL11 cable that goes to IN1). The digital phase rotation for usual POY Yarm locking is 106 degrees, so the analog POY11_Q channel contains most of the signal. I then set the IN1 gain of the servo to 0dB, and looked at the CM_Slow signal - I changed the whitening gain of this channel to +18dB (to match that used for POY11_I and POY11_Q), and found that I had to apply a digital gain of 0.5 to get the PDH horns in the usual POY11_I signal and the CM_Slow signal to line up. There was also a sign inversion. Then I was able to use the digital LSC system and lock the Y arm cavity length to the PSL frequency by actuating on ETMY, using CM_Slow as an error signal. A comparison of the in-loop POY11_I ASD when the arm is locked is shown in Attachment #1 - CMslow seems to be dominated by some kind of electroncis noise above ~100 Hz, so possibly needs more whitening (even though the nominal whitening filter is engaged)?

Anyway, now that I have this part of the servo working, the next step is to try and engage the AO path and achieve a higher BW lock of the Y arm cavity to the PSL frequency (= IMC length). Maybe it makes more sense to actuate on MC2 for the slow path.

Attachment 1: YARM_CMslow.pdf
YARM_CMslow.pdf
  15041   Wed Nov 20 21:29:28 2019 gautamUpdateLSCPRG ~13

After the QPD fix, both arms report consistent buildup - see Attachment #1. The peak values touch ~250, corresponding to a PRG of ~13. The IFO becomes critically coupled at PRG=15. I am finding that the 3f signal offsets are changing as a function of the CARM offset, and this could be responsible for the lock breaking as I approach 0 CARM offset. I found that I could maintain a more stable and deterministic transition to zero CARM offset by dynamically adjusting the 3f PRCL error signal offset to keep the REFL11 signal approximately at 0. Some shaking seems to have commenced so I am breaking for now.

Note that I find scattered throughout the elog references to a similar problem of the PRMI losing lock as the CARM offset is reduced, e.g. here. But haven't stumbled across what the resolution was, the PRFPMI could be locked pretty easily in 2015 I remember.

Attachment 1: PRG13.pdf
PRG13.pdf
  15040   Wed Nov 20 17:52:00 2019 gautamUpdateLSCQPD MEDM screen update

Koji and I had noticed that there was some discrepancy between the switchable gain stages of the EX and EY QPDs. Sadly, there was no indication that these switches even exist on the QPD MEDM screen. Yehonathan and I rectified this today. Both EX and EY Transmon QPDs now have some extra info (see Attachment #1). We physically verified the indicated quadrant mapping for the EX QPD (see previous elog in this thread for the details), and I edited the screen accordingly. EY QPD still has to be checked. Note also that there is an ND=0.4 + ND=0.2 filter and some kind of 1064nm light filter installed in series on the EX QPD. The ND filters have a net transmission T~25%.

After making the EX and EY QPDs have the same switchable gain settings (I also reset the trans normalization gains), the angular motion witnessed is much more consistent between the two now - see Attachment #2. The high-frequency noise of the sum channel is somewhat higher for EX than EY - maybe the ND filters are different on the two ends?

Note that there was an extra factor of 40 gain on the EX QPD relative to EY during the lock yesterday. So really, the signals were probably just getting saturated. Now that the gains are consistent between the ends, it'll be interesting to see how balanced the buildup of the two arms is. There still remains the problem that the MICH loop was too unstable, which probably led to to excess arm power fluctuations.

There is a mark on the QPD surface that is probably dirt (since we never have such high power transmitted through the ETM to damage the QPD). I'll try cleaning it up at the next opportunity.

Attachment 1: newLookQPD.png
newLookQPD.png
Attachment 2: TRX_TRY_comparison.pdf
TRX_TRY_comparison.pdf
Attachment 3: IMG_8186.JPG
IMG_8186.JPG
  15039   Wed Nov 20 17:20:24 2019 YehonathanUpdateLSCQPD Investigation

{Gautam, Yehonathan}

In search of the source of discrepancy between the QPD readings in the X and Y arms, we look into the schematics of the QPD amplifier - DCC #D990272.

We find that there are 4 gain switches with the following gain characteristics (The 40m QPD whitening board has an additional gain of 4.5):

S4 S3 S2 S1 V/A
0 0 0 0 2e4
0 0 0 1 2e5
0 0 1 0 4e4
0 0 1 1 4e5
0 1 0 0 1e5
0 1 0 1 1e6
0 1 1 0 2e5
0 1 1 1 2e6
1     0 5e2
1     1 5e3

Switch 4 bypasses the amps controlled by switch 2 and 3 when it is set to 1 so they don't matter in this state.

Note that according to elog-13965 the switches are controlled through the QPD whitening board by a XT1111a Acromag whose normal state is 1.

Also, according to the QPD amplifier schematics, the resistor on the transimpedance, controlled by switch 1, is 25kOhm. However, according to the EPICS it is actually 5kOhm. We verify this by shining the QPD with uniform light from a flashlight and switching switch1 on and off while measuring the voltages of the different segments. The schematics should be updated on the DCC.

Surprisingly, QPDX switches where 0,0,0,0 while QPDY switches where 1,0,0,1. This explains the difference in their responses.

We check by shining a laser pointer with a known power on the different segments of QPDX that we get the expected number of counts on the ADC and that the response of the different segments is equal.
 

gautam edits:

  1. Lest there be confusion, the states of the switches in the (S1, S2, S3, S4) order are (0,0,0,0) for QPDX and (0,1,0,1) for QPDY.
  2. The Acromag XT1111 is a sinking BIO unit - so when the EPICS channel is zero, the output impedance is low and the DUT (i.e. MAX333) is shorted to ground. So, the state of the MAX333 shown on the schematics corresponds to EPICS logic level 1, and the switched state corresponds to logic level 0.
  3. For the laser pointer test, we used a red laser pointer. Using a power meter, we measured ~100uW of 632nm power. However, we think this particular laser pointer had failing batteries or something because the spot looked sometimes brighter/dimmer to the eye. Anyways, we saw ~10,000 ADC counts when illuminating a single segment (with the QPD gain switches at the 0,0,0,0 setting, before we changed anything). We expect 100uW * 0.4 A/W * 500 V/A * 10 * 40 * 4.5 * 3267.8 cts/V = ~12000 cts. So everything seems to check out. We changed the gain to the 5kohm setting and bypassed the subsequent gain stages, and saw the expected response too. The segments were only balanced to ~10%, but presumably this can be adjusted by tweaking digital gains.
  15038   Wed Nov 20 12:14:17 2019 gautamUpdateLSCLocking - progress

I took a look at the TRX/TRY RIN reported in the single arm POX/POY lock, and compared the performance of the two available PDs at each of the two ends. Attachment #1 shows the results. Some remarks:

  1. The noise performance of the two QPDs at each end isn't identical - is there some transimpedance gain difference?
  2. The lower plot shows the angular motion reported by each QPD when the arm cavity is locked. The EX QPD seems much more sensitive than the EY QPD.
  3. I estimate that in this condition, each photodiode is receiving ~20uW of power, corresponding to a shot noise limited RIN of ~10^-7. None of the photodiodes saturate this bound.
  4. There are some ND filters placed in front of the QPDs at both ends. Do we really need these ND filters? I estimate that for the highest buildups, we will have ~10kW * 15ppm * 0.5 ~75mW of power incident on the QPD, so ~20mW per segment. Assuming silicon responsivity of 0.2 A/W, a transimpedance of 1kohm would give us 4V of signal. But the schematic shows higher transimpedance. Do we still have the switching capability for this QPD?
Quote:

Next steps:

  1. Check the EX QPD / TRX situation.
Attachment 1: TRX_TRY_comparison.pdf
TRX_TRY_comparison.pdf
  15037   Wed Nov 20 01:07:18 2019 gautamUpdateLSCLocking - progress

Summary:

  1. CARM offset was reduced to 0 with the PRMI locked.
  2. TRY levels touched ~200 (Recycling gain ~10, IFO is still undercoupled).
  3. TRX level never went so high - I suspect QPD issues or clipping in the beampath.

Details:

  • Attachment #1 is a StripTool summary of the lock - encouragingly, the PRMI stayed locked for several 10s of minutes even when the CARM offset was brought to 0.
  • The MICH signal was pretty glitchy - we increased the gain of the MICH and PRCL loops and thought we saw some improvement, but needs more quantitative investigation.
  • Main differences in locking procedure today were:
    • Added some POPDC to the MICH/PRCL trigger elements in addition to POP22
    • Tried adding a DARM offset before doing the final steps of CARM offset reduction, and then zerod the DARM offset too.
  • The TRX level never went as high as TRY - even though on the CRT monitors, both arms seemed to saturate somewhat more evenly. Potentially the EX QPD needs a checkout, or there is some clipping in the in-air TRX path. Although, puzzilingly, the POXDC level never goes as high as POYDC either. So maybe the buildup is really lower in the XARM? For the daytime tomorrow.

Next steps:

  1. Check the EX QPD / TRX situation.
  2. Figure out how to make the PRMI lock more stable as I reduce the CARM offset.
  3. Start figuring out the CM board, as we'd want to do the handoff to RF at some point.
Attachment 1: PRFPMI.png
PRFPMI.png
  15036   Tue Nov 19 21:53:57 2019 gautamUpdatePEMFollow-up on seismometer discussion

The shaking started earlier today than yesterday, at ~9pm local time.

While the IFO is shaking, I thought (as Jan Harms suggested) I'd take a look at the cross-spectra between our seismometer channels at the dominant excitation frequency, which is ~1.135 Hz. Attachment #1 shows the phase of the cross spectrum taken for 10 averages (with 30mHz resolution) during the time period when the shaking was strong yesterday (~1500 seconds with 50% overlap). The logic is that we can use the relative phasing between the seismometer channels to estimate the direction of arrival and hence, the source location. However, I already see some inconsistencies - for example, the relative phase between BS_Z and EX_Z suggests that the signal arrives at the EX seismometer first. But the phasing between EX_Y and BS_Y suggests the opposite. So maybe my thinking about the problem as 3 co-located sensors measuring plane-wave disturbances originating from the same place is too simplistic? Moreover, Koji points out that for two sensors separated by ~40m, for a ground wave velocity of 1.5 km/s, the maximum phase delay we should see between sensors is 30 msec, which corresponds to ~10 degrees of phase. I guess we have to undo the effects of the phasing in the electronics chain.

Does anyone have some code that's already attempted something similar that I can put the data through? I'd like to not get sucked into writing fresh code.

🤞 this means that the shaking is over for today and I get a few hours of locking time later today evening.

Another observarion is that even after the main 1.14 Hz peak dies out, there is elevated seismic acitivity reported by the 1-3 Hz BLRMS band. This unfortunately coincides with some stack resonance, and so the arm cavity transmission reports greater RIN even after the main peak dies out. Today, it seems that all the BLRMS return to their "nominal" nighttime levels ~10 mins after the main 1.14 Hz peak dies out.

Attachment 1: seisxSpec.pdf
seisxSpec.pdf
  15035   Tue Nov 19 15:08:48 2019 gautamUpdateCDSVertex models rebooted

Jon and I were surveying the CDS situation so that he can prepare a report for discussion with Rolf/Rich about our upcoming BHD upgrade. In our poking around, we must have bumped something somewhere because the c1ioo machine went offline, and consequently, took all the vertex models out. I rebooted everything with the reboot script, everything seems to have come back smoothly. I took this opportunity to install some saturation counters for the arm servos, as we have for the CARM/DARM loops, because I want to use these for a watch script that catches when the ALS loses lock and shuts stuff off before kicking optics around needlessly. See Attachment #1 for my changes.

Attachment 1: armSat.png
armSat.png
  15034   Mon Nov 18 21:04:38 2019 gautamUpdateLSCLocking - some ideas

Some ideas Koji suggested:

  1. Try approaching the CARM offset zero point "from the other side" - i.e. start with a CARM offset of the opposite sign (I typically use negative CARM offset).
  2. With the PRMI locked, try bringing one arm onto resonance while the other arm is held off resonance. 

For the second idea, it is convenient to be able to control the arms in the XARM/YARM basis as opposed to the CARM/DARM basis as we usually do when going through the locking sequence. After some fiddling, I am able to reliably execute this transition, and achieve a state where the FP arm cavities are resonant for the IR with the ALS beat note frequency being the error signal being used. Some important differences:

  1. The frequency actuator (ETM) is weaker in this case than in the CARM/DARM basis (where MC2 is the frequency actuator) due to the longer length of the arm cavity (and for ETMX, the higher coil driver series resistance). I had to twiddle the limits of the servo banks to accommodate this. 
  2. The ALS error signal is significantly noisier than POX/POY. Hence, the control signal RMS is often in danger of saturating the DAC range. I implemented a partial fix by adding a 1st order Butterworth LPF with 1kHz corner frequency. According to the model, this eats <5 degrees of phase at the desired UGF of ~150 Hz.
  15033   Mon Nov 18 16:32:15 2019 gautamUpdateComputersZITA: started upgrade from Ubuntu 14 LTS to 18 LTS

the upgrade seems to have been successfully executed - the machine was restarted at ~430pm local time. Projector remains off and diagnostic striptools are on the samsung.

Quote:

and so it begins...until this is finished I have turned off the projector and moved the striptools to the big TV (time to look for Black Friday deals to replace the projector with a 120 inch LED TV)

  15032   Mon Nov 18 14:32:53 2019 gautamUpdatePEMFollow-up on seismometer discussion

The nightly seismic activity enhancement continued during the weekend. It always shows up around 10pm local time, persists for ~1 hour, and then goes away. This isn't a show stopper as long as it stops at some point, but it is annoying that it is eating up >1 hour of possible locking time. I walked over to CES, no one there admitted to anything - there is an "Earth Surface Dynamics Laboratory" there that runs some heavy equipment right next to us, but they claim they aren't running anything post ~530pm. Rick (building manager ?) also doesn't know of anything that turns on with the periodicity we see. He suggested contacting Watson but I have no idea who to talk to there who has an overview of what goes on in the building. 😢 

  15031   Fri Nov 15 18:59:08 2019 ranaUpdateComputersZITA: started upgrade from Ubuntu 14 LTS to 18 LTS

and so it begins...until this is finished I have turned off the projector and moved the striptools to the big TV (time to look for Black Friday deals to replace the projector with a 120 inch LED TV)

  15030   Fri Nov 15 12:16:48 2019 gautamUpdatePEMFollow-up on seismometer discussion

Attachment #1 is a spectrogram of the BS sesimometer signals for a ~24 hour period (from Wednesday night to Thursday night local time, zipped because its a large file). I've marked the nearly pure tones that show up for some time and then turn off. We need to get to the bottom of this and ideally stop it from happening at night because it is eating ~1 hour of lockable time.

We considered if we could look at the phasing between the vertex and end seismometers to localize the source of the disturbance.

Attachment 1: BS_ZspecGram.pdf.zip
  15029   Fri Nov 15 12:08:04 2019 gautamUpdateLSCPOPDC whitening board

The DC port of the Bias-Tee is routed to (a modified version of) the iLIGO whitening board. This has the well-known problem of the protection diodes of the LT1125 quad-op-amp lowering the (ideally infinite) input impedance of the first gain stage (+24 dB). To be sure as to how much signal we can put into this port (in anticipation of trying some variable finesse PRFPMI locking but also for general book-keeping), I tested the usable input range by driving a triangle wave at ~3 Hz and changing the amplitude of the signal until we observed saturation. We found that we could drive a 10 Vpp signal at which point there was evidence of some clipping (it was asymmetric, the top end of the signal was getting clipped at +14,000 cts while the bottom end still looked like a triangle wave at -16,000 counts). Anyway we probably don't want to exceed +/- 10,000 counts on this channel. This is consistent with Hartmut's statement of having +/- 4V of usable range (although the counts he mentions are twice what I saw yesterday).

Other discussion points between Rana, Koji and Gautam:

  1. Conside putting an in-vacuum (Silicon ?) QPD for the PRC angular motion sensing
    • In-vacuum will yield lower acoustic noise coupling
    • Bring the photocurrent out and do the transimpedance amplification in air 
    • Use a large area QPD so as to be more tolerant to alignment drifts without having to introduce picomotors (but how much does the POP spot actually drift and is this feasible?)
  2. Is there some better telescope configuration for the existing in-air QPD?
    • What is the correct Gouy-phase for this to be able to best sense the PRC cavity axis motion?
  15028   Fri Nov 15 11:58:12 2019 gautamUpdateLSCoff the bad Violin filters

The clue was that the loop X arm POX loop looked to have low (<3dB)) gain margin around 600 Hz (and again at 700 Hz). Attachment #1 shows a comparison of the OLTF for this loop (measured using the IN1/IN2 method) before and after our change. We hypothesize that one of the violin filters that were turned off had non-unity DC gain, because I had to lower the loop gain by 20% after these turn-offs to have the same UGF. I updated the snap files called by the arm locking scripts.

I think I caught all the places where the FM settings are saved, but some locking scripts may still try and turn on some of these filters, so let's keep an eye on it.

Quote:

We turned off many excessive violin mode bandstop filters in the LSC.

Attachment 1: violinFix.pdf
violinFix.pdf
Attachment 2: newViolinConfig.png
newViolinConfig.png
  15027   Fri Nov 15 00:18:41 2019 ranaUpdatePEMFollow-up on seismometer discussion

The large ground motion at 1 Hz started up again tonight at around 23:30. I walked around the lab and nearby buildings with a flashlight and couldn't find anything whumping. The noise is very sinusoidal and seems like it must be a 1 Hz motor rather than any natural disturbance or traffic, etc. Suspect that it is a pump in the nearby CES building which is waking up and running to fill up some liquid level. Will check out in the morning.

Estimate of displacement noise based on the observed MC_F channel showing a 25 MHz peak-peak excursion for the laser:

dL = 25e6 * (13 m / (c / lambda)

      = 1 micron

So this is a lot. Probably our pendulum is amplifying the ground motion by 10x, so I suspect a ground noise of ~0.1 micron peak-peak.

(this is a native PDF export using qtgrace rather than XMgrace. uninstall xmgrace and symlink to qtgrace.)

Attachment 1: MCshake.pdf
MCshake.pdf
  15026   Thu Nov 14 23:56:18 2019 ranaUpdateLSCoff the bad Violin filters

We turned off many excessive violin mode bandstop filters in the LSC.

Due to some feedforward work by Jenne or EQ some years ago, we have had ~10 violin notches on in the LSC between the output matrix and the outputs to the SUS.

They were eating phase, computation time, and making ~3 dB gain peaking in places where we can't afford it. I have turned them off and Gautam SDF safed it.

Offensive BS shown in brown and cooler BS shown in blue.

To rotate the DTT landscape plot to not be sideways, use this command (note that the string is 1east, not least):
pdftk in.pdf cat 1east output out.pdf

Attachment 1: out.pdf
out.pdf
  15025   Thu Nov 14 12:11:04 2019 ranaUpdatePEMFollow-up on seismometer discussion

at 1 Hz' this effect is not large so that's real translation. at lower frequencies a ground tilt couples to the horizontal sensors at first order and so the apparent signal is amplified by the double integral. drawing a free body diagram u can c that

x_apparent = (g / s^2) * theta

but for vortical this not tru because it already measures the full free fall and the tilt only shows up at 2nd order

  15024   Wed Nov 13 23:40:15 2019 gautamUpdatePEMFollow-up on seismometer discussion

Here is some disturbance in the spacetime curvature, where the local gradient of the metric seems to have been modulated (in the "downward" as well as in the other two orthogonal Cartesian directions) at ~1 Hz - seems real as far as I can tell, all the suspensions were being shaken about and all the seismometers witnessed it, though the peak is pretty narrow. A broader, less prominent peak also shows up around 0.5 Hz. We couldn't identify any clear source (no LN2 fill-up / obvious CES activity). This event lasted for ~45 mins, and stopped around 2315 local time. Shortly (~5min) after the ~1 Hz peak died down, however, the 3-10 Hz BLRMS channel reports an increase by ~factor of 2. 

Onto trying some locking now that the suspensions have settled down somewhat.

Quote:

this is due to the Equivalence Principle: local accelerations are indistinguishable from spacetime curvature. On a spherical Earth, the local gradient of the metric points in the direction towards the center of the Earth, which is colloquially known as "down".

Attachment 1: seisAll_20191111_1Hz.pdf
seisAll_20191111_1Hz.pdf
  15023   Wed Nov 13 20:15:56 2019 ranaUpdatePEMFollow-up on seismometer discussion

this is due to the Equivalence Principle: local accelerations are indistinguishable from spacetime curvature. On a spherical Earth, the local gradient of the metric points in the direction towards the center of the Earth, which is colloquially known as "down".

Quote:

I don't understand why the z-axis motion reported by the T240 is ~10x lower at 10 mHz compared to the X and Y motions. Is this some electronics noise artefact?

 

  15022   Wed Nov 13 19:34:45 2019 gautamUpdatePEMFollow-up on seismometer discussion

Attachment #1 shows the spectra of our three available seismometers over a period of ~10ksec.

  • I don't understand why the z-axis motion reported by the T240 is ~10x lower at 10 mHz compared to the X and Y motions. Is this some electronics noise artefact?
  • The difference in the low frequency (<100mHz) shapes of the T240 compared to the Guralps is presumably due to the difference in the internal preamps / readout boxes (?). I haven't checked yet.
  • There is almost certainly some issue with the EX Guralp. IIRC this is the one that had cabling issues in the past, and also is the one that was being futzed around for Tctrl, but also could be that its masses need re-centering, since it is EX_X that is showing the anomalous behaviour.
  • The coherence structure between the other pairs of sensors is consistent.

Attachment #2 shows the result of applying frequency domain Wiener filter subtraction to the POP QPD (target) with the vertex seismometer signals as witness channels.

  • The dataset was PRMI locked with the carrier resonant, ETMs misaligned.
  • The dashed lines in these plots correspond to the RMS for the solid line with the same color.
  • For both PIT and YAW, I am using BS_X and BS_Y seismometer channels for the MISO filter inputs.
  • In particular for PIT, I notice that I am unable to get the same level of performance as in the past, particularly around ~2-3 Hz.
  • The BS seismometer health indicators don't signal any obvious problems with the seismometer itself - so something has changed w.r.t. how the ground motion propagates to the PR2/PR3? Or has the seismometer sensing truly degraded? I don't think the dataset I collected was particularly bad compared to the past, and I confirmed similar performance with a separate PRMI lock from a different time period.
Attachment 1: seisAll_20191111.pdf
seisAll_20191111.pdf
Attachment 2: ffPotential.pdf
ffPotential.pdf
  15021   Thu Nov 7 17:55:37 2019 shrutiUpdateComputer Scripts / ProgramsPython packages on donatella

Today I realized that pip and other python2,3 packages were installed in the conda base environment, so after running

conda activate

I could run the python-GPIB scripts to interface with the Agilent.

Although, I did have to add a python2 kernel to jupyter/ipython, which I did in a separate conda environment:

conda create -n ipykernel_py2 python=2 ipykernel
source activate ipykernel_py2
python -m ipykernel install --user
Quote:

I've installed pyepics on Donatella running

sudo yum install pyepics

Pip and ipython did not seem to be installed yet.

 

  15020   Thu Nov 7 17:46:10 2019 shrutiUpdateALSAM measurement at X end

Some details:

- There was a SR560+SR785 (not connected for measurement) placed near the X end which I moved; it is now behind the electronics rack by the X arm beam tube (~15m away).

- Also, for the AM measurement I moved the AG5395A from behind the PSL setup to the X end, where it now is.

- By toggling the XGREEN shutter, I noticed that the cavity was not resonant before I disconnected anything from the setup since the spot shape kept changing, but I proceeded anyway. 

- Because Rana said that it was important for me to mention: the ~5 USD blue-yellow crocs (that I now use) work fine for me.

The AM Measurement:

1. The cables were calibrated with the DC block in the A port (for a A/R measurement)

2. The cable to the PZT was disconnected from the pomona box and connected to the RF out of the NA, the PD output labelled 'GREEN_REFL' was also disconnected and connected to the B port via a DC block. 

3. The ITMX was 'misaligned'. (This allowed the reflected green PD output as seen on the oscilloscope to stabilize.)

4. The PZT is modulated in frequency and the residual amplitude modulation (as observed in the measured reflected green light) is plotted, ref. Attachment 1. The parameters for the plotted data in the attachment were:

# AG4395A Measurement - Timestamp: Nov 07 2019 - 17:04:07
#---------- Measurement Parameters ------------
# Start Frequency (Hz): 10000.0, 10000.0
# Stop Frequency (Hz): 10000000.0, 10000000.0
# Frequency Points: 801, 801
# Measurement Format: LOGM, PHAS
# Measuremed Input: AR, AR
#---------- Analyzer Settings ----------
# Number of Averages: 8
# Auto Bandwidth: On, On
# IF Bandwidth: 300.0, 300.0
# Input Attenuators (R,A,B): 0dB 10dB 20dB 
# Excitation amplitude = -10.0dBm

 

 

------------------------------------

Update (19:13 7thNov19):  When the ITMX was intentionally misaligned, Rana and I checked to see if the Oplevs were turned off and they were. But while I was casually checking the Oplevs again, they were on! 

Not sure what to do about this or what caused it. 

Quote:

[Shruti, Rana]

- At the X end, we set up the network analyzer to begin measurement of the AM transfer function by actuation of the laser PZT.

- The lid of the PDH optics setup was removed to make some checks and then replaced.

- From the PDH servo electronics setup the 'GREEN_REFL' and 'TO AUX-X LASER PZT' cables were removed for the measurement and then re-attached after.

- The signal today was too low to make a real measurement of the AM transfer function, but the GPIB scripts and interfacing was tested. 

 

Attachment 1: AMTF20191107.png
AMTF20191107.png
  15019   Wed Nov 6 20:34:28 2019 KojiUpdateIOOPower combiner loss (EOM resonant box installed)

Gautam and I were talking about some modulation and demodulation and wondered what is the power combining situation for the triple resonant EOM installed 8 years ago. And we noticed that the current setup has additional ~5dB loss associated with the 3-to-1 power combiner. (Figure a)

N-to-1 broadband power combiners have an intrinsic loss of 10 log10(N). You can think about a reciprocal process (power splitting) (Figure b). The 2W input coming to the 2-port power splitter gives us two 1W outputs. The opposite process is power combining as shown in Figure c. This case, the two identical signals are the constructively added in the combiner, but the output is not 20Vpk but 14Vpk. Considering thge linearity, when one of the port is terminated, the output is going to be a half. So we expect 27dBm output for a 30dBm input (Figure d). This fact is frequently oversight particularly when one combines the signals at multiple frequencies (Figrue e). We can avoid this kind of loss by using a frequency-dependent power combiner like a diplexer or a triplexer.

Attachment 1: power_combiner.pdf
power_combiner.pdf
  Draft   Wed Nov 6 20:34:08 2019 KojiUpdateIOOEOM resonant box installed

 

Quote:

[Mirko / Kiwamu]

 The resonant box has been installed together with a 3 dB attenuator.

The demodulation phase of the MC lock was readjusted and the MC is now happily locked.

 

(Background)

We needed more modulation depth on each modulation frequency and so for the reason we installed the resonant box to amplify the signal levels.

Since the resonant box isn't impedance matched well, the box creates some amount of the RF reflections (#5339).

In order to reduce somewhat of the RF reflection we decided to put a 3 dB attenuator in between the generation box and the resonant box.

 

(what we did)

 + attached the resonant box directly to the EOM input with a short SMA connector.

 + put stacked black plates underneath the resonant box to support the wight of the box and to relief the strain on the cable between the EOM and the box.

 + put a 3 dB attenuator just after the RF power combiner to reduce RF reflections.

 + readjusted the demodulation phase of the MC lock.

 

(Adjustment of MC demodulation phase)

 The demodulation phase was readjusted by adding more cable length in the local oscillator line.

After some iterations an additional cable length of about 30 cm was inserted to maximize the Q-phase signal.

So for the MC lock we are using the Q signal, which is the same as it had been before.

 

 Before the installation of the resonant box, the amplitude of the MC PDH signal was measured in the demodulation board's monitor pins.

The amplitude was about 500 mV in peak-peak (see the attached pictures of the I-Q projection in an oscilloscope). Then after the installation the amplitude decreased to 400 mV in peak-peak.

Therefore the amplitude of the PDH signal decreased by 20 %, which is not as bad as I expected since the previous measurement indicated 40 % reduction (#2586).

 

 

  15017   Wed Nov 6 19:26:57 2019 gautamUpdatePSLSome PSL cable admin

Koji and I taked about cleaning up some of the flaky cable situation on the PSL table a while ago. The changes were implemented and are documented in Attachment #1. Now the Pomona box between the Thorlabs HV Driver and the NPRO head is sitting on the PSL table (sandwiched between some teflon pieces I found in cabinet S4 along the south arm), and the cables between these two devices are better strain relieved. I turned off the Thorlabs HV supply while working on the PMC table. The IMC could be locked after this work. Probably won't solve the long standing FSS mysteries but probably can't hurt.

Unrelated to this work: I also removed a Bias tee that was just hanging out on top of the FSS electronics, which was used for the modeSpec project.

Attachment 1: PSLcableAdmin.jpg
PSLcableAdmin.jpg
  15016   Wed Nov 6 17:45:34 2019 gautamUpdateLSC~

Here is a comparison of the response of various DoFs in our various RFPD sensors for two different CARM offsets. Even in the case of the smaller CARM offset of ~1kHz, we are several linewidths away from the resonance. Need to do some finesse modeling to make any meaningful statement about this - why is the CARM response in REFL11 apparently smaller for the smaller CARM offset?

If you mistrust my signal processing, the GPS times for which I ran the sensing lines are:

CARM offset = ~30kHz (arm transmission <0.02) --- 1257064777+5min

CARM offset = ~1kHz (arm transmission ~5) --- 1257065566+5min

Quote:

Summary:

There seems to be stronger-than-expected coupling between CARM and the 3f sensors. 

  15015   Wed Nov 6 17:05:45 2019 gautamUpdateLSCCARM calibration

Summary:

A coarse calibration of the CARM error point (when on ALS control) is 7.040 +/- 0.030 kHz/ct. This corresponds to approximately 0.95nm/ct. I typically lose the PRMI lock when the CARM offset is ~0.2 cts, which means I am about 1kHz away from the resonance. This is >10 CARM linewidths.

Details:

The calibration was done by sweeping the CARM offset (no PRM) and identifying the arm cavity FSRs by looking for peaks in TRX / TRY. Attachment #1 shows the scan, while Attachment #2 shows a linear fit to the FSRs. In Attachment #2, the frequency axis is taken from the phase tracker servo, which was calibrated by injecting a "known" frequency with the Marconi, and there is good agreement to the expected FSR with 37.79 m long arm cavities. There is much more info in the scan (e.g. modulation depths, mode matching to the arm cavities etc) which I will extract later, but if anyone wants the data (pre-downsampled by me to have a managable filesize), it's attached as a .zip file in Attachment #3.

Attachment 1: CARMscan.pdf
CARMscan.pdf
Attachment 2: CARMcalib.pdf
CARMcalib.pdf
Attachment 3: scan.hdf5.zip
  15014   Wed Nov 6 02:08:48 2019 gautamUpdateLSCLocking updates

Summary:

There seems to be stronger-than-expected coupling between CARM and the 3f sensors. 

Details:

Full analysis tomorrow, but I collected sensing matrix measurements with lines driven in PRCL,MICH and CARM at a couple of CARM offsets. I also wanted to calibrate the CARM offset to physical units so I ran some scans of the CARM offset and collected the data so I can use the arm cavity FSR to calibrate CARM. Koji suggested using REFL165_I for PRCL and REFL165_Q for MICH control - this would allow us to see if the problem was with the 1f sideband only. While the lock could be established, we still couldn't push the arm powers above 10 without breaking the PRMI lock. While changing the CARM offset, we saw a significant shift in the DC offset level of the out-of-loop REFL33_I signal. Need to think about what this means...

  15013   Tue Nov 5 12:37:50 2019 gautamUpdatePEMT240 interface unit pulled out

I removed the Trillium T240 DAQ interface unit from 1X4 for investigation.

It was returned to the electronics rack and all the connections were re-made. Some details:

  1. The board is indeed a D1000749-v2 as Koji said it is. There is just an additional board (labelled D1001872 but for which there is no schematic on the DCC) inside the 1U box that breaks out the D37 connector of the v2 into 3 D15 connectors. I took photos.
  2. Armed with the new cable Chub got, and following the manual, I ran the re-centering routine.
    • Now all the mass-monitoring position voltages are <0.3 V DC, as the manual tells me they should be.
    • I noticed that when the seismometer is just plugged in and powered, it takes a few minutes for the mass monitoring voltages to acquire their steady state values.
    • The V indicator reported ~-2V DC, and the W indicator reported -3.9V DC.
    • While running the re-centering routine, I monitored the mass-position indicator voltages (via the backplane D15 connector) on an oscilloscope. See Attachment #1 for the time series. The data was rather noisy, I don't know why this is, so I plot the raw data in light colors and a filtered version in darker colors. Also, there seems to be a gain of x2 in the voltages on the backplane relative to what the T240 manual tells me I should expect, and the values reported when I query the unit via the serial port.
    • We should ideally just install another Acromag ADC in the c1susaux box and acquire these and other available diagnostic information, since the signals are available.
    • We should also probably check the mass position indicator values in a few days to see if they've drifted off again.
    • Looking at the raw time series / spectra of the BS channels, I see no obvious signatures of any change. 
    • I will run a test by locking the PRC and looking for coherence between the seismometer data and angular motion witnessed by the POP QPD, as this was what signalled my investigation in the first place.

Update 445pm: Seems to have done something good - the old feedforward filters reduce the YAW RMS motion by a factor of a few. Pitch performance is not so good, maybe the filter needs re-training, but I see coherence, see Attachment #2 for the frequency domain WF.

Attachment 1: T240_recenter.pdf
T240_recenter.pdf
Attachment 2: ffPotential.pdf
ffPotential.pdf
  15012   Tue Nov 5 11:52:27 2019 gautamUpdateLSCLocking notes

Summary:

I am still unable to achieve arm powers greater than TRX/TRY ~10 while keeping the PRMI locked. A couple of times, I was able to get TRY ~50, but TRX stayed at ~10, or even dropped a little, suggestive of a DARM offset? On the positive side, the ALS system seems to work pretty reliably, and I can keep the arms controlled by ALS for several tens of minutes.

Details:

  • Despite my POP beam path improvements, I saw the POP22 level drop as I lowered the CARM offset.
  • One strange feature last night was that with the arms held off resonance using ALS, I had to flip the sign and increase the gain by ~x2 of the REFL33_I-->PRCL loop in order to lock the PRMI. This was confirmed by locking on the 1f error signals and measuring the ratio of the response between the 1f and 3f signals while shaking PRCL using DTT swept sine.
  • At different CARM offsets, I noted that the DC offset level on the 1f photodiodes (i.e. REFL11 and AS55) were changing significantly.
  • I ran a measurement of the sensing matrix with the arm powers hovering around ~10, which is just before I lose the PRMI lock - managed to stay locked for >5 minutes, but the sensing matrix seems to suggest that the REFL33 demod angle needs to be rotated - maybe this is the reason why the PDH horn-to-horn voltage of REFL33 is lower now than it was last week? No idea why that should be, I was around the LSC rack but if the situation is so fragile, seems hopeless.
  • MICH sensed by REFL165_Q still seems stable, so that's good...
  • So my best hypothesis at the moment is that the PRCL optical gain is falling as I reduce the CARM offset (due to DC offset? or something else?). Needs some detailed modeling for more insight, I'm out of ideas for tests to run while locking as I've gone through the full gamut of OLTF and sensing matrix measurements at various CARM offsets without getting any clues as to what's going on.
Attachment 1: PRMI3f_ALS_Nov4sensMat.pdf
PRMI3f_ALS_Nov4sensMat.pdf
  15011   Mon Nov 4 19:02:25 2019 YehonathanUpdatePSLMapping the PSL electronics

I created a spreadsheet (Attached) by taking Koji's c1psl sheet from slow_channel_list and filtering out the channels that do not need an Acromag. I added in the QPD channels that are relevant to the PSL from the c1iool0 sheet.

I began mapping the PSL related Eurocrates connectors to their respective VME channels starting with the PMC electronics.

I am confused about the TTFSS interface (D040423): While it is a Eurocrate card, in the schematics it seems to have 50 pin connectors.

I found old wiring schematics that might help with identifying the channels once the connector issue is clarified.

 

 

Attachment 1: PSL_Wirings_-_Sheet1_(1).pdf
PSL_Wirings_-_Sheet1_(1).pdf
  15010   Mon Nov 4 16:06:58 2019 gautamUpdateLSCPOP optical path

I did some re-alignment of the POP beam on the IX in air table. Here are the details:

  1. Attachment #1 - optical layout.
  2. With the PRC locked with the carrier resonant (no arm cavities), there is ~300uW of DC power incident on the Thorlabs PDA10CF, which serves as POP22, POP110 and POPDC photosensor.
    • See this elog for the signal paths.
    • On a scope, this corresponded to ~1.8 V DC of voltage. This is in good agreement with the expected transimpedance gain of 10 kOhms and responsivity of ~0.65 A/W given on the datasheet.
    • This is also in agreement with the ~6000 ADC counts I see in the CDS system (although there are large fluctuations). 
  3. These was significant misalignment of the beam on this photodiode at some point:
    • Previously, I had used the CDS system to walk the beam on thde photodiode to try and maximize the power.
    • Today I took a different approach - triggered the MICH and PRCL loops on REFLDC (instead of the usual POPDC / POP22) so I could freely block the beam.
    • I found that there is a fast (f=35mm) lens to make the beam small enough for the PDA10CF. The beam was somewhat mis-centered on this strongly curved optic, and I suspect it was amplifying small misalignments. Anyway it is much better centered now (see Attachment #2) and I have a much stronger POPDC signal (by a factor of ~2-3, see Attachment #3).
    • The ASS dither alignment now shows much more consistent behavior - minimizing REFLDC maximises POPDC, see Attachment #4.
    • I took this opportunity to take some spectra/time-series of the PD output with the interferometer in this configuration. 

Tangentially related to this work - I took the nuclear option and did a hard reboot of the c1susaux Acromag crate on Sunday to fix the EPICS issue - it seems to be gone for now, see Attachment #5.

Attachment 1: IMG_8027.JPG
IMG_8027.JPG
Attachment 2: lensRealignment.jpg
lensRealignment.jpg
Attachment 3: POPrealigned.png
POPrealigned.png
Attachment 4: POPdither.png
POPdither.png
Attachment 5: PRMfixed.png
PRMfixed.png
  15009   Mon Nov 4 15:29:47 2019 gautamUpdateLSCPOP signal path

There are many versions of the POP22 signal path I found on the elog, e.g. this thread. But what I saw at the LSC rack was not quite in agreement with any of those. So here is the latest greatest version.

Since the 2f signals are mainly indicators of power buildups and are used for triggering various PDH loops, I don't know how critical some of these things are, but here are some remarks:

  1. There is no Tee + 50 ohm terminator after the minicircuits filters, whose impedance in the stopband are High-Z (I have been told but never personally verified).
  2. The RF amplifier used is a Minicircuits ZFL-1000-LN+. This has a gain of 20dB and 1dB compression output power spec of 3dBm. So to be safe, we want to have not more than -20dBm of signal at the input. On a 50-ohm scope (AC coupled), I saw a signal that has ~100mVpp amplitude (there is a mixture of many frequencies so this is not the Vpp of a pure sinusoid). This corresponds to -16dBm. Might be cutting it a bit close even after accounting for cable loss and insertion loss of the bias tee.
  3. We use a resistive power splitter to divide the power between the POP22 and POP110 paths, which automatically throws away 50% of the RF power. A better option is the ZAPD-2-252-S+.
  4. The Thorlabs PDA10CF photodiode (not this particular one) has been modelled to have a response that can be approximated by a complex pole pair with Q=1 at ~130 MHz. But we are also using this PD for measuring the 110 MHz PD which is a bit close to the band edge?
Attachment 1: POPchain.pdf
POPchain.pdf
  15008   Mon Nov 4 13:26:04 2019 YehonathanUpdatePSLUp to date sketch of the 1x1 and 1x2 Eurocrates

Done.

Quote:

Thanks. Please update this wiki page too.

https://wiki-40m.ligo.caltech.edu/Electronics/ElectronicsRacks#A1X1

  15007   Mon Nov 4 11:41:28 2019 shrutiUpdateComputer Scripts / ProgramsEpics installed on donatella

I've installed pyepics on Donatella running

sudo yum install pyepics

Pip and ipython did not seem to be installed yet.

  15006   Sat Nov 2 17:08:34 2019 YehonathanUpdatePSLUp to date sketch of the 1x1 and 1x2 Eurocrates

Thanks. Please update this wiki page too.

https://wiki-40m.ligo.caltech.edu/Electronics/ElectronicsRacks#A1X1

  15005   Sat Nov 2 16:36:55 2019 YehonathanUpdatePSLUp to date sketch of the 1x1 and 1x2 Eurocrates

I reproduced Gautam's sketch of the 1x1 and 1x2 Eurocrates into a pdf image that contains links to the appropriate DCCs in the legend (see attachement).

Attachment 1: 1x1_1X2_Eurocrates_with_links.pdf
1x1_1X2_Eurocrates_with_links.pdf
  15004   Thu Oct 31 10:44:40 2019 gautamUpdatePSLPMC re-locked

PMC got unlocked at ~4am. I re-locked it. Also tweaked the input pointing into the cavity. The misalignment was mostly in pitch.

There was also a loud buzzing in the control room due to the audio cable being improperly seated in the mixer. I re-seated it.

  15003   Wed Oct 30 23:12:27 2019 KojiUpdateSUSPRM suspension issues

Sigh... hard loch

  15002   Wed Oct 30 19:20:27 2019 gautamUpdateSUSPRM suspension issues

While I was trying to lock the PRMI this evening, I noticed that I couldn't move the REFL beamspot on the CCD field of view by adjusting the slow bias voltages to the PRM. Other suspensions controlled by c1susaux seem to respond okay so at first glance it isn't a problem with the Acromag. Looking at the OSEM sensor input levels, I noticed that UL is much lower than the others - see Attachment #1, seems to have happened ~100 days ago. I plugged the tester box in to check if the problem is with the electronics or if this is an actual shorting of some pins on the physical OSEM as we had in the past. So PRM watchdog is shutdown for now and there is no control of the optic available as the cables are detached. I will replace the connections later in the evening.

Update 10pm:

  1. Measured coil inductances with breakout board and LCR meter - all 5 coils returned ~3.28-3.32 mH.
  2. Measured coil resistances with breakout board and DMM - all 5 coils returned ~16-17 ohms.
  3. Checked OSEM PD capacitance (with no bias voltage) using the LCR meter - each PD returned ~1nF.
  4. Checked resistance between LED Cathode and Anode for all 5 LEDs using DMM - each returned Hi-Z.
  5. Checked resistance between PD Cathode and Anode for all 5 PDs using DMM - each returned ~430 kohms.
  6. Checked that I could change the slow bias voltages and see a response at the expected pins (with the suspension disconnected).

Since I couldn't find anything wrong, I plugged the suspension back in - and voila, the suspect UL PD voltage level came back to a level consistent with the others! See Attachment #2.

Anyway, I had some hours of data with the tester box plugged in - see Attachment #3 for a comparison of the shadow sensor readout with the tester box (all black traces) vs with the suspension plugged in, local damping loops active (coloured traces). The sensing noise re-injection will depend on the specifics of the  local damping loop shapes but I suspect it will limit feedforward subtraction possibilities at low frequencies.

However, I continue to have problems aligning the optic using the slow bias sliders (but the fast ones work just fine) - problem seems to be EPICS related. In Attachment #4, I show that even though I change the soft PITCH bias voltage adjust channel for the PRM, the linked channels which control the actual voltages to the coils take several seconds to show any response, and do so asynchronously. I tried restarting the modbus process on c1susaux, but the problem persists. Perhaps it needs a reboot of the computer and/or the acromag chassis? I note that the same problem exists for the BS and PRM suspensions, but not for ITMX or ITMY (didn't check the IMC optics). Perhaps a particular Acromag DAC unit is faulty / has issues with the internal subnet?

Attachment 1: PRMUL.pdf
PRMUL.pdf
Attachment 2: PRMnormal.pdf
PRMnormal.pdf
Attachment 3: PRM-Sensors_noise.pdf
PRM-Sensors_noise.pdf PRM-Sensors_noise.pdf
Attachment 4: PRMsuspensionWonky.png
PRMsuspensionWonky.png
ELOG V3.1.3-